Fermi National Laboratory


Accelerator Update for July 4, 2005 - July 6, 2005

Monday July 4
The day shift began with the Main Injector (MI) off for MI-20 transformer repairs, with Operations monitoring the Recycler stash and the Antiproton Source (Pbar) stack, and with MiniBooNE taking beam.

Operations prepared the Collider Detector at Fermilab (CDF) for an access at 8:18 AM. They came out of access five hours later.

A Booster RF station (BRF13) tripped off on a ground fault indication for the first of many times over the next three hours. It reset okay each time.

At 9:11 AM, a Pbar Debuncher device (D:IB) tripped off. Operators reset it, but then discovered a LCW leak coming from DRF1-5. They called in an expert.

At 12:15 PM, BRF13 tripped off a bunch more times. Operators left the station off.

Operations put the TeV through a dry squeeze at 1:53 PM.

At 3:10 PM, Operations reported that the Tevatron Electron Lens (TEL) would not turn on. An operator notified an expert and the expert had the TEL back on by 4:23 PM.

Operations reported at 9:06 PM that technicians were conducting the final tests on the MI-20 transformer.

Operations reported at 10:28 PM that they had established beam to the MI dump.

At 11:25 PM, a NuMI power supply (E:V118) tripped off due to a bad capacitor. Also, a Lambertson (I:LAM62) tripped off and wouldn't reset. Operators contacted an expert who said that the NuMI techs had been here all day and had to go home. NuMI would not run. He suggested two other people to investigate LAM62. They were called in.

Tuesday July 5
The midnight shift began with Operations monitoring the stack and stash, with techs working on the LAM62 power supply, and with MiniBooNE taking beam.

The techs completed their LAM62 repair at 2:10 AM.

Operations put the TeV into shot setup at 2:13 AM.

Operations established combination store 4250 at 6:09 AM with an initial luminosity of 98.4E30.

Pbar resumed stacking at 6:29 AM.

At 6:45 AM, the Meson MCenter experimenters reported that they couldn’t control their analyzing magnet. Operators discovered a CAMAC problem and informed the Controls experts.

The Controls expert replaced a bad Switchyard CAMAC power supply at 8:12 AM.

Operations had beam established to NuMI and MCenter by 11:23 AM.

Operations put the Recycler into shot setup to stash antiprotons at 12:36 PM.

BRF13 was brought back on line at 3:38 PM.

A Pbar expert began a two-hour study period at 8:37 PM.

Wednesday July 6
The midnight shift began with Operations monitoring store 4250 and the stash, with Pbar stacking, and with the following experiments taking beam: MiniBooNE, NuMI, and MCenter.

At 12:51 AM, MCenter experimenters reported a problem with their SWICs. Operators informed experts who investigated and then reverted to some older code. Operators rebooted the front end and the SWIC readbacks resumed at 2:30 AM.

The Plan for Wednesday
The plans for today are to end the store this morning and then let CDF and D-Zero access their collision halls. The next shot setup will begin around noon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/5/2005   email Fermilab